In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 21 vowels and 15 consonants, what is the probability you will choose a consonant then a vowel.
21 vowels.....15 consonants.......added = 36 tiles
probability of first picking a consonant : 15/36 reduces to 5/12
replacing it...
probability that second pick is a vowel : 21/36
probability of both : 5/12 * 21/36 = 105 / 432 reduces to 35/144 <===
The probability that choosing a consonant then a vowel from the bag after replacement is 35/144.
"Probability means possibility. It is a branch of mathematics that deals with the occurrence of a random event. The value is expressed from zero to one".
For the given situation,
The bag consists of letters 21 vowels and 15 consonants.
Total number of letters = 36
The event is the probability that choosing a consonant then a vowel from the bag after replacement.
Probability of choosing a consonant = \(\frac{15}{36}\)
Probability of choosing a vowel = \(\frac{21}{36}\)
Thus the probability of the event,
\(P(e) = \frac{15}{36}\) × \(\frac{21}{36}\)
⇒ \(P(e)=\frac{35}{144}\)
Hence we can conclude that the probability that choosing a consonant then a vowel from the bag after replacement is 35/144.
Learn more about probability here
brainly.com/question/13604758
#SPJ2